OpenStudy (anonymous):

Lt x tends to infinity {(x+6)/(x+1)}^{x+4}

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \left( \left( x+6 \right) \div \left( x+1 \right)\right)^{x+4}\]

OpenStudy (anonymous):

You guys want the answer, I just need the working

OpenStudy (mani_jha):

is the answer 1?

OpenStudy (anonymous):

No its \[e ^{5}\]

OpenStudy (mani_jha):

I think differentiating this with respect to x might give that. Let me work on it

OpenStudy (anonymous):

You need a teorem in order to do this actually or else you are just wasting time